Google next year will have FCFTE of $29.10, that are expected to grow at 25% for the next four years. After that, Google’s cash flows will grow only at about 5.5% forever. If the cost of equity capital for Google is 10.20%, and Google has 700 million shares outstanding, what should be the share price for Google today
Answer #1

The share price for Google = Present value of the future FCFTE/Number of shares outstanding

The present value of the future FCFTE

PV = 29.10/1.1020^1 + (29.10 * 1.25)/1.1020^2 + (29.10 * 1.25^2)/1.1020^3 + (29.10 * 1.25^3)/1.1020^4 + (29.10 * 1.25^4)/1.1020^5 + (29.10 * 1.0505)/(0.1020 - 0.055) * 1/1.1020^5

PV = 26.4065335753 + 29.9529645818 + 33.9756857779 + 38.5386635412 + 43.714455015 + 400.2056496429

PV = $572.793952134 million

The share price for Google = 572.793952134/700

The share price for Google = $0.8182770745

Can you please upvote? Thank You :-)